satellite73 (satellite73):

find the first 4 terms \[a_1=\frac{3}{2}; a_{n+1}=\frac{n^2+1}{n(a_n)}\]

OpenStudy (anonymous):

i got my answer but apparently i was wrong this from @lopez_hatesmath

OpenStudy (anonymous):

maybe some fresh eyes would help

OpenStudy (anonymous):

nvm sorry to bother you

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

swagg

OpenStudy (anonymous):

it was this \[a_1=\frac{3}{2}; a_{n+1}=\frac{n^2+1}{n}\times a_n\]

OpenStudy (anonymous):

that is, the \(a_n\) was in the NUMERATOR

OpenStudy (anonymous):

so now it is not so bad replace \(n=1\) on the right hand side to get \(a_2\)

OpenStudy (anonymous):

you get \[a_2=\frac{1^2+1}{1}\times \frac{3}{2}\] \[a_2=2\times\frac{3}{2}\] \[a_2=3\]

OpenStudy (anonymous):

now replace \(n\) by 2 on the right hand side to get \[a_3=\frac{2^2+1}{2}\times 3\] \[a_3=\frac{5}{2}\times 3\] \[a_3=\frac{15}{2}\]

OpenStudy (anonymous):

how are we doing so far?

OpenStudy (anonymous):

good :)

OpenStudy (anonymous):

one more? \[a_4=\frac{3^2+1}{3}\times \frac{15}{2}\] \[a_4=\frac{10}{3}\times \frac{15}{2}\] \[a_4=25\]

OpenStudy (anonymous):

okayy i got it!

OpenStudy (anonymous):

not so bad that is the idea, i thought the term was in the denominator which is why i was screwing it up

OpenStudy (anonymous):

gotcha thanks for the help man.

OpenStudy (anonymous):

yw good luck with the next one, but it works the same so you should be good to go